OpenStudy (anonymous):

Using complete sentences, explain how to graph y = -1/5x - 2 using the slope-intercept method.

OpenStudy (anonymous):

hey angelina inbox me.. ill help u with this problem also

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

ok hold on im doing it right now:)

OpenStudy (anonymous):

Plot (0,-2). Then, move a unit down and 5 units right. Plot that point and connect both pts with the line...

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

should be.. ill lok more into it if u need me to

OpenStudy (anonymous):

oh yea and talk to me on inbow got to aks u sooemthing

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how did you get (0,2)?

OpenStudy (anonymous):

i said plot (0,-2). Then, move a unit down and 5 units right. Plot that point and connect both pts with the line...

OpenStudy (anonymous):

then u will get th answer

OpenStudy (anonymous):

what is the answer to graph the equation IN COMPLETE SENTENCES

OpenStudy (anonymous):

ok wait then!!!!!!>:(

OpenStudy (anonymous):

i been waiting

OpenStudy (anonymous):

ok u want me to do the problem u cant do soo... u have to wait till im done

OpenStudy (anonymous):

ok and that what im doing waiting

OpenStudy (anonymous):

Given line is in slope intercept form y = mx + c slope = m = -1/5 y intercept = c =-2

OpenStudy (anonymous):

its pretty much straight forward from there

OpenStudy (anonymous):

at least figure out the complete sentence could be

OpenStudy (anonymous):

haha fine

OpenStudy (anonymous):

ok kool since u want to act like that to me even tho i did nothing to u

OpenStudy (anonymous):

u dont have to act like ur my friend if u arent

OpenStudy (anonymous):

what u talking bout.... you be acting crazy i an't did nuttin to you

OpenStudy (anonymous):

no today since u told me that u had a boyfriend u been acting weird for some reason

OpenStudy (anonymous):

naw this is how i act i mean am i have attitude hell i am Italian

OpenStudy (anonymous):

so that gives u an excus to be mad at somebody who did nothing to u

OpenStudy (anonymous):

i aint mad BOFL

OpenStudy (anonymous):

wat does BOFL mean

OpenStudy (anonymous):

Balling On Floor Laughing

OpenStudy (anonymous):

oh inbox me real wuick gotta tell u something again that cant be put in the open

OpenStudy (anonymous):

i did